OpenStudy (anonymous):

Solve 2x – 5 = 3(2x + 5).. can u please show me the steps? so i can get it.. thanks

OpenStudy (anonymous):

2x - 5 = 6x + 15 -20 = 4x -5 = x

OpenStudy (katrinakaif):

First distribute: 2x - 5 = 6x + 15 Now bring all the x' to one side -5 = 6x - 4x + 15 -5 = 2x + 15 -20 = 2x x = -10

OpenStudy (anonymous):

2x-5=6x+15 -5=6x-2x +15 -5-15=6x-2x -20=4x -5 = x

OpenStudy (anonymous):

katrina that was good but i saw a little mistake when u moved the 2x to the other side.. but thank you because u made me think and that was good.. thanks for your help